Rylee88:

For which of the following inequalities is x => 4 a solution? x + 3 => 1 3x= > 1 x - 3= > 1

Rylee88:

=> is the sign with A less than, bigger than, with the line under,

IhelpUuHELPme:

This ones' fun >.> \[x \ge 4\] This is similar to a one step equation What we have to do is SOLVE the inequality to see if it gives the final answer as x => 4 So \[x + 3 \ge 1\] We subtract 3 from both sides

IhelpUuHELPme:

Now we simplify (: \[x \ge -2\]

IhelpUuHELPme:

This inequality didn't give us the value of x => 4 So we can eliminate this answer choice

IhelpUuHELPme:

So far, do you get what Im saying?

IhelpUuHELPme:

We're doing Inverse operations to where we can get x by itself

Rylee88:

Yeah sorry lol

IhelpUuHELPme:

\[3x \ge 1\] Inverse operations >.> 3x means multiplication meaning we divide 3 on both sides which gives you the inequality : \[x \ge \frac{ 1 }{ 3}\]

IhelpUuHELPme:

Which now leaves you with \[x - 3 \ge 1\] Keyword: INVERSE meaning OPPOSITE - 3 is a negative number you can also say it's subtraction, opposite of subtractuon is addition You add 3 on both sides 3 and 3 cancel out which leaves x by itself 3 + 1 = 4 Now we simplify \[x \ge 4\] The inequality: \[x - 3 \ge 1\] When we do inverse operations, the solution is \[x \ge 4\] Which is what the question is asking for

IhelpUuHELPme:

Your answer is The inequality \[x - 3 \ge 1\] give you the solution \[x \ge 4\]

Rylee88:

Ohhhhhhhhhhhhhh ty

IhelpUuHELPme:

Do you understand tho? xD

Rylee88:

Kindaaaaaaaa

IhelpUuHELPme:

And you're welcome

Rylee88:

x≥4 this means? lol

IhelpUuHELPme:

Kinda- Lemme *simplify* it- math puns smh When having an inequality, you find the solution by using inverse operations For example the inequality x - 3 => 1 To get x by itself, we have to "remove" the number 3 We add 3 on both sides, we add because of inverse operations When we simplify we get x => 4 (:

IhelpUuHELPme:

@rylee88 wrote:
x≥4 this means? lol
It means the number that will be in place of the variable x is equal to or greater than 4

Rylee88:

Thats the equation...

IhelpUuHELPme:

if the sign faced towards the 4 It would mean the number that will be in place of the variable x is equal to or less than 4

IhelpUuHELPme:

@rylee88 wrote:
Thats the equation...
No no no- It's the solution for the inequality x - 3 => 1

Rylee88:

OHHHHHHHhhh

IhelpUuHELPme:

YESS XD

IhelpUuHELPme:

When we remove the greater than sign we get x = 4 But in this problem, It's what we call an inequality, meaning we have a greater than or less than sign included

IhelpUuHELPme:

Does that make more sense?

Rylee88:

Yes

IhelpUuHELPme:

Glad to help!

Rylee88:

We can do the rest tomorrow xD

IhelpUuHELPme:

Ight xd

Rylee88:

Thank you so much you don't know how much this means to me lol

IhelpUuHELPme:

You're welcome and anytime xd